# TCJA-SNN: Temporal-Channel Joint Attention for Spiking Neural Networks

### v2

@@ IV-A1 Dataset Introduction
  • DVS128 Gesture The DVS128 Gesture [36] dataset is an event-stream dataset composed of 11 kinds of hand gestures from 29 subjects under three different illumination conditions, directly captured with the DVS128 camera.
@@ IV-A2 Neuromorphic Dataset Preprocessing
where $x_{i}$ and $y_{i}$ event’s coordinate, $p_{i}$ represents the event. In order to reduce computational consumption, we group events into $T$ slices, where $T$ is the network’s time simulation step. A frame in the integrated Frame data, denoted as $F(j)$, refers to the pixel value at position $(p,x,y)$, represented as $F(j,p,x,y)$. It is obtained by integrating events indexed between $j_{l}$ and $j_{r}$ from the event stream. The process can be described as:
@@ IV-B Network Architecture
The network architectures designed for the different datasets are listed in Tab. I.
@@
Hyperparameter CIFAR10-DVS N-Caltech 101 DVS128 Fashion-MNIST
Learning Rate $31e-$ 3 $31e-$ 3 $31e-$ 3 $31e-$ 3
Batch Size 64 32 16 128
$T$ 10 14 20 8
@@ IV-C Network Implementation
We train and test our method on a workstation equipped with two Tesla P4 and two Tesla P10 GPUs. As the memory consumption, we use the Tesla P10 to train and test the CIFAR10-DVS dataset, N-Caltech 101 dataset, and DVS128 Gesture dataset, and use the Tesla P4 to train and test the Fashion-MNIST dataset. For different datasets, hyperparameters are listed in Tab. II. We set the learning rate to $31e-$ 3 in each dataset empirically, and train the same epochs as [23] except N-Caltech 101, which is not tested in [23]. we enable the automatic mixed precision in N-Caltech 101 and DVS128 Gesture for the excessive resolution ($180\times 240$ and $128\times 128$). We also detach the reset process in backpropagation, which is validated to enhance performance.
@@ IV-D Comparison with Existing Classification SOTA Works
The performance of two TCJA-SNN variants is compared with some SOTA models in Tab. III and Tab. IV. We train and test two variants with SpikingJelly [41] package based on PyTorch [42] framework, resulting in enhanced performance across all tasks. Some studies [16, 43, 12] substitute binary spikes with floating-point spikes in whole or in part and retain the same temporal forward pipeline as SNN to obtain improved classification accuracy. Thus, we devise two variants to validate the efficiency of TCJA-SNN by utilizing the TET loss function. On CIFAR10-DVS, we obtain 3.4% advantage over the prior method with binary spikes. On N-Caltech 101, with only 14 time steps, we get a 15.7% increase over the prior best work. On DVS128, we get an accuracy of 99.0%, which is higher than TA-SNN [12] using three times fewer simulation time steps. Furthermore, by using a basic 7-layer CNN on the static dataset Fashion MNIST, our method can achieve the highest classification accuracy with the fewest simulation time steps. Overall, with binary spikes, TCJA-SNN simulates no-more time steps while getting higher performance. Furthermore, our method can achieve higher classification accuracy by adopting the non-binary spike technique.
@@
Method Binary Spikes CIFAR10-DVS N-Caltech 101 DVS128 $T$ Step Acc. $T$ Step Acc. $T$ Step Acc. SLAYER [44]NeurIPS-2018 - - - - 1600 93.4 HATS [45]CVPR-2018 N/A N/A 52.4 N/A 64.2 - - DART [46]TPAMI-2019 N/A N/A 65.8 N/A 66.8 - - NeuNorm [11]AAAI-2019 230-292 60.5 - - - - Rollout [38]Front. Neurosci-2020 48 66.8 - - 240 97.2 DECOLLE [47]Front. Neurosci-2020 - - - - 500 95.5 LIAF-Net [43]TNNLS-2021 10 70.4 - - 60 97.6 tdBN [9]AAAI-2021 10 67.8 - - 40 96.9 PLIF [23]ICCV-2021 20 74.8 - - 20 97.6 TA-SNN [12]ICCV-2021 10 72.0 - - 60 98.6 SEW-ResNet [16]NeurIPS-2021 16 74.4 - - 16 97.9 Dspike [31]NeurIPS-2021 10 $*{}^{*}$ SALT [48]Neural Netw-2021 20 67.1 20 55.0 - - TET [29]ICLR-2022 10 $*{}^{*}$ DSR [33]CVPR-2022 10 $*{}^{*}$ TCJA-SNN 10 $*{}^{*}$ 14 78.5 20 99.0 TCJA-TET-SNN 10 $*{}^{*}$ 14 82.5 20 98.2 $*{}^{*}$ With Data Augmentation.
TABLE IV: Static Fashion-MNIST accuracy.
Method Binary Spike Time Step Accuracy ST-RSBP [49]NIPS-2019 400 90.1 LISNN [50]IJCAI-2020 20 92.1 PLIF [23]ICCV-2021 8 94.4 TCJA-SNN 8 94.8 TCJA-TET-SNN 8 94.6
TABLE V: Accuracy of different blocks.
Figure 6: Convergence of compared SNN methods on DVS128 Gesture.
@@ IV-E Comparison with Existing Image Generation Works
In this experiment, we build a fully spiking variation autoencoder (FSVAE) for image generation. Moreover, we replace the original image decoding way with our novel method by calculating the average output on the temporal dimension after the TCJA block. The workflow chart of this FSVAE with TCJA applied on image decoding is shown in Fig. 8. And, we apply log-likelihood evidence lower bound (ELBO) as the loss function:
@@
We employ the AdamW optimizer [51], which trains 300 epochs at 0.001 learning rate and 0.001 weight decay. The batch size is set to 256. Moreover, the time step is set to 16. The performance of the TCJA image decode is compared with some SOTA models in Tab. VI. Because this method can make full use of the powerful temporal information, the Inception Score (IS) shows SOTA results compared to the original FSVAE [52] and the same structure ANN. And our TCJA image decoding outperforms better on all metrics for CIFAR10 datasets. Moreover, results on CelebA and MNIST are further visualized in Fig. 9 - 10, which demonstrates that our generated images are visually better than the previous method [52].
@@ IV-F Ablation Study
To thoroughly examine the impact of the TLA and CLA modules, we conducted a series of ablation studies. The results, as presented in Tab. V, indicate that the CLA module plays a crucial role in enhancing performance. This can be attributed to the fact that, in most SNN designs, the number of simulation time steps is significantly fewer than the number of channels. Consequently, the CLA module is able to extract additional relevant features compared to the TLA module. Furthermore, it is worth noting that the TCJA module consistently outperformed other models across all tested datasets. This outcome underscores the effectiveness of the CCF layer incorporated within the TCJA module, further reinforcing its potential for achieving superior performance.
@@ IV-G1 Kernel Size
To verify the effectiveness of our proposed CCF mechanism, we devise a variant method that substitutes addition for multiplication of $\mathcal{T}_{i,j}$ and $\mathcal{C}_{i,j}$ in the Eq.  6. The results are shown in Tab. VII.
TABLE VII: Test accuracy on three datasets with different CCF operations.
@@ IV-G3 Convergence
We also empirically demonstrate the convergence of our proposed method, as shown in Fig. 6. Specifically, Fig. 6 illustrates the performance trend of vanilla LIF-SNN, PLIF [23] and our proposed TCJA-SNN for 1000 epochs. As the training epoch increases, the performance trend of our proposed method becomes more stable and converges to a higher level. Moreover, the TCJA-SNN can achieve the SOTA performance when only training about 260 epochs, which demonstrates the efficacy of the proposed TCJA.
@@ IV-G5 Theoretical Analysis on Receptive Field
Recalling Eq.  4 and Eq.  5, through two 1-D convolutions along different dimensions, we construct two CCS in a vertical relationship, which are stored in $\mathcal{T}$ and $\mathcal{C}$. In particular, TCJA is to construct a CCS, which can perceive a larger area while realizing feature interaction in different directions. This cross-receptive field is able to abolish the limitations caused by the monotonic dimension, thus bringing performance improvements to the network.
@@ IV-G7 Energy Consumption Analysis
Compared to the Analog Neural Network (ANN), the SNN consumes less energy due to its sparser firing and poorer processing accuracy. Owing to the binary spikes, each operation in SNN consists of a single floating-point (FP) addition. In ANN, on the other hand, each operation computes a dot product as a multiply-accumulate (MAC) calculation consisting of one floating-point (FP) multiplication and one FP addition. Consequently, SNNs use less energy than ANNs in the same network design. This discrepancy can also be validated in 45nm CMOS technology, where the energy cost of each SNN operation is 5.1$\times$ lower than that of each 32-bit ANN MAC operation (0.9pJ vs. 4.6pJ) [53], allowing us to examine the energy consumption of each network architecture.
We assess the energy consumption in the network for classifying the DVS128 dataset with both ANN and SNN. first, we assess the spiking rate, floating-point operations per second (FLOPS) of each layer, the result is shown in Tab. VIII.
@@
For ANN, we could calculate the energy consumption by $FLOPS\times MAC\;energy\;cost$; for SNN, the energy cost should be quantified by $FLOPS\times SNN\;operation\;energy\;cost\times spiking\;rate$. The final power consumption calculation results are $1.90\times 10^{-3}$ (TCJA-SNN) and $10.00\times 10^{-3}$ (ANN), where our TCJA-SNN costs 5.26×5.26$\times$5.26 × lower energy consumption compared to its ANN version.
@@ IV-G8 Propagation Pattern
Both the temporal and the spatial domains are involved in SNN forward propagation. In intuition, the forward propagation computation graph of SNN may be arranged in a step-by-step pattern, which is illustrated in Fig. 12. Step-by-step means we evaluate the output of the entire network at the first timestep, with updates to the hidden state of the spiking neurons, and then evaluate the subsequent timesteps in the same way. Besides, the layer-by-layer pattern is also extensively used, which entails performing a spatial forward propagation procedure in which we calculate the output of the first layer at all timesteps as the input of the second layer, then retrieve the output of the last layer at all timesteps. Unlike parallel computing environments like GPU, where the layer-by-layer pattern is preferred, neuromorphic devices operate more like a step-by-step pattern. It can be proved that the output of the network in the two patterns is mathematically equivalent.
@@ V Conclusion
In this paper, we propose the TCJA mechanism, which innovatively recalibrates temporal and channel information in SNN. Specifically, instead of utilizing a generic fully connected network, we use 1-D convolution to build the correlation between frames, reducing the computation and improving model performance. Moreover, we propose a CCF mechanism to realize joint feature interaction between temporal and channel information. Sufficient experiments verify the effectiveness of our method with SOTA results on four datasets, i.e., CIFAR10-DVS (83.3%), N-Caltech101 (82.5%), DVS128 (99.0%) and Fashion-MNIST (94.8%). In addition to its outstanding performance in classification tasks, TCJA-SNN also exhibits a competitive performance in image generation tasks. To the best of our knowledge, this study represents the pioneering application of the SNN-attention mechanism to both high-level classification and low-level generation tasks. Remarkably, our approach has achieved state-of-the-art performance in both domains, thus making a significant advancement in the field. However, the insertion of TCJA still resulted in a relatively sizable boost in the number of parameters. In future work, we believe this method can easily be integrated into the neuromorphic chip for the hardware-friendly 1-D convolution operation and the binary spiking network structure.

### v3

@@ IV-A1 Dataset Introduction
  • DVS128 Gesture The DVS128 Gesture [45] dataset is an event-stream dataset composed of 11 kinds of hand gestures from 29 subjects under three different illumination conditions, directly captured with the DVS128 camera. In this paper, we employs all 11 gesture categories for the purpose of classification.
@@
  • CIFAR10/100 The CIFAR10/100 dataset [47] consists of 60,000 32 $\times$ 32 images with 3 channels in 10/100 classes. There are 50,000 training images and 10,000 testing images.
@@ IV-A2 Neuromorphic Dataset Preprocessing
where $x_{i}$ and $y_{i}$ event’s coordinate, $p_{i}$ represents the event. In order to reduce computational consumption, we group events into $T$ slices, where $T$ is the network’s time simulation step. A frame in the integrated frame data, denoted as $F(j)$, refers to the pixel value at position $(p,x,y)$, represented as $F(j,p,x,y)$. It is obtained by integrating events indexed between $j_{l}$ and $j_{r}$ from the event stream, where Here, $j_{l}$ represents the initial timestamp for accumulation, and $j_{r}$ denotes the final timestamp. The process can be described as:
@@ IV-B Network Architecture
The architectures of networks corresponding to various datasets are enumerated in Table I. In the construction of each network, Kaiming [50] initialization is methodically applied to both convolutional and fully-connected layers.
@@
Hyperparameter CIFAR10-DVS N-Caltech 101 DVS128 Fashion-MNIST CIFAR10 CIFAR100
Optimizer Adam Adam Adam Adam SGD SGD
Learning Rate $31e-$ 3 $31e-$ 3 $31e-$ 3 $31e-$ 3 $11e-$ 1 $11e-$ 1
Batch Size 64 32 16 128 128 128
$T$ 10 14 20 8 6/4 6/4
@@
For the CIFAR10/100 dataset, we employ the MS-ResNet architecture, as detailed in Ref. [52], to validate the effectiveness of the TCJA on deep residual neural networks. Specifically, we utilize the standard MS-ResNet-18 architecture for classifying the CIFAR datasets. The TCJA module is integrated at the bottom of each MS-ResNet block.
@@ IV-C Network Implementation
We train and test our method on a workstation equipped with two Tesla P4 and two Tesla P10 GPUs. As the memory consumption, we use the Tesla P10 to train and test the CIFAR10-DVS dataset, N-Caltech 101 dataset, and DVS128 Gesture dataset, and use the Tesla P4 to train and test the Fashion-MNIST, CIFAR10/100 dataset. In the various datasets under consideration, the hyperparameters are detailed in Table II. The learning rate has been empirically set to $1\times 10^{-3}$ for each dataset when utilizing the Adam optimizer. Conversely, for the implementation involving ResNet with the SGD optimizer, a higher learning rate of $1\times 10^{-1}$ has been employed, as the SGD optimizer necessitates a more substantial learning rate. and train the same epochs as [30] except N-Caltech 101, which is not tested in [30]. we enable the automatic mixed precision in N-Caltech 101 and DVS128 Gesture for the excessive resolution ($180\times 240$ and $128\times 128$). We strategically detach the reset process during backpropagation, a technique increasingly recognized for its effectiveness in optimizing SNNs. The detach operation decouples the reset operation from the computational graph. Such a detachment has been empirically validated to enhance network performance, offering a more efficient approach to managing the dynamics of SNNs [21, 30, 2]. This process ensures that the essential learning dynamics are retained while unnecessary computational complexities are minimized, thereby improving the overall efficacy of the network.
@@ IV-D Comparison with Existing Classification SOTA Works
The performance of two TCJA-SNN variants is compared with some SOTA models in Tab. III, Tab. IV and Tab. V. We train and test two variants with SpikingJelly [53] package based on PyTorch [54] framework, resulting in enhanced performance across all tasks. Some studies [21, 55, 14] substitute binary spikes with floating-point spikes in whole or in part and retain the same temporal forward pipeline as SNN to obtain improved classification accuracy. Thus, we devise two variants to validate the efficiency of TCJA-SNN by utilizing the TET loss function. On CIFAR10-DVS, we obtain 1.7% advantage over the prior method with binary spikes. on the N-Caltech 101 dataset, we achieved a classification accuracy of 82.5%, surpassing previous work by 1.6%. On DVS128, we get an accuracy of 99.0%, which is higher than TA-SNN [14] using three times fewer simulation time steps. Furthermore, by using a basic 7-layer CNN on the static dataset Fashion MNIST, our method can achieve the highest classification accuracy with the fewest simulation time steps. In the context of the CIFAR10 and CIFAR100 datasets, the implementation of TCJA demonstrates a significant improvement over the baseline models [52] that do not incorporate TCJA. Specifically, there is an enhancement of 2.08% and 3.83% in classification accuracy for CIFAR10 and CIFAR100, respectively. Additionally, our method surpasses the current SOTA models, as referenced in Ref. [56], by margins of 0.84% for CIFAR10 and 0.93% for CIFAR100. Overall, with binary spikes, TCJA-SNN simulates no-more time steps while getting higher performance. Furthermore, our method can achieve higher classification accuracy by adopting the non-binary spike technique.
@@
Method Binary Spikes CIFAR10-DVS N-Caltech 101 DVS128 $T$ Step Acc. $T$ Step Acc. $T$ Step Acc. SLAYER [57]NeurIPS-2018 - - - - 1600 93.4 HATS [58]CVPR-2018 N/A N/A 52.4 N/A 64.2 - - DART [59]TPAMI-2019 N/A N/A 65.8 N/A 66.8 - - NeuNorm [13]AAAI-2019 230-292 60.5 - - - - Rollout [48]Front. Neurosci-2020 48 66.8 - - 240 97.2 DECOLLE [60]Front. Neurosci-2020 - - - - 500 95.5 LIAF-Net [55]TNNLS-2021 10 70.4 - - 60 97.6 tdBN [11]AAAI-2021 10 67.8 - - 40 96.9 PLIF [30]ICCV-2021 20 74.8 - - 20 97.6 TA-SNN [14]ICCV-2021 10 72.0 - - 60 98.6 SEW-ResNet [21]NeurIPS-2021 16 74.4 - - 16 97.9 Dspike [40]NeurIPS-2021 10 75.4 - - - - SALT [61]Neural Netw-2021 20 67.1 20 55.0 - - TET [36]ICLR-2022 10 83.2 - - - - DSR [42]CVPR-2022 10 77.3 - - - - Event Transformer [62] N/A 71.2 N/A 78.9 - - STCA-SNN [63]Front. Neurosci-2023 10 81.6 14 80.9 - - TCJA-SNN 10 80.7 14 78.5 20 99.0 TCJA-TET-SNN 10 83.3 14 82.5 20 98.2 With Data Augmentation.
TABLE IV: The comparison between the proposed methods and existing SOTA techniques on static CIFAR datasets. (Bold: the best)
Methods Architecture CIFAR10 CIFAR100 $T$ Step Acc. $T$ Step Acc. ANN2SNN [64]AAAI-2023 ResNet-18/ResNet-20 32 95.42 32 65.50 tdBN [11]AAAI-2021 Spiking-ResNet-19 6 93.16 6 71.12 TET [36]ICLR-2022 Spiking-ResNet-19 6 94.50 6 74.72 RecDis [65]CVPR-2022 Spiking-ResNet-19 6 94.71 6 74.10 GLIF [56]NeurIPS-2022 Spiking-ResNet-19 6 95.03 6 77.35 MS-ResNet [52] MS-ResNet-18 6 93.79 6 74.45 TCJA-SNN MS-ResNet-18 6 95.87 6 78.28 MS-ResNet-18 4 95.60 4 77.72 ANN [52] MS-ResNet-18 N/A 96.41 N/A 80.67
TABLE V: Static Fashion-MNIST accuracy.
Method Binary Spike Time Step Accuracy ST-RSBP [66]NeurIPS-2019 400 90.1 LISNN [67]IJCAI-2020 20 92.1 PLIF [30]ICCV-2021 8 94.4 TCJA-SNN 8 94.8 TCJA-TET-SNN 8 94.6
@@ IV-E Comparison with Existing Image Generation Works
In this experiment, we build a fully spiking variation autoencoder (FSVAE) for image generation with TCJA. Moreover, we replace the original image decoding way with our novel method by calculating the average output on the temporal dimension after the TCJA block. The workflow chart of this FSVAE with TCJA applied on image decoding is shown in Fig. 6. And, we apply log-likelihood evidence lower bound (ELBO) as the loss function:
@@
We employ the AdamW optimizer [68] for image generating tasks, which trains 300 epochs at 0.001 learning rate and 0.001 weight decay. The batch size is set to 256. Moreover, the time step is set to 16. The performance of the TCJA image decode is compared with some SOTA models in Tab. VI. Because this method can make full use of the powerful temporal information, the Inception Score (IS) shows SOTA results compared to the original FSVAE [69] and the same structure ANN. And our TCJA image decoding outperforms better on all metrics for CIFAR10 datasets. Moreover, results on CelebA and MNIST are further visualized in Fig. 7 - 8, which demonstrates that our generated images are visually better than the previous method [69]. In addition to the primary evaluations, our model was also compared with other methods based on the Spiking VAEs, such as image decoding based on temporal attention (TAID) [70] and Efficient Spiking VAE (ESVAE) [71], which have been proposed recently. Despite these methods being specifically designed for low-level image reconstruction tasks, our approach remains competitive. It demonstrates robust performance, reflecting its adaptability and effectiveness in comparison to these specialized models.
@@ IV-F Ablation Study
Figure 9: Convergence of compared SNN methods on DVS128 Gesture.
TABLE VII: Accuracy of different blocks.
To thoroughly examine the impact of the TLA and CLA modules, we conducted a series of ablation studies. The results, as presented in Tab. VII, indicate that the CLA module plays a crucial role in enhancing performance. This can be attributed to the fact that, in most SNN designs, the number of simulation time steps is significantly fewer than the number of channels. Consequently, the CLA module is able to extract additional relevant features compared to the TLA module. Furthermore, it is worth noting that the TCJA module consistently outperformed other models across all tested datasets. This outcome underscores the effectiveness of the CCF layer incorporated within the TCJA module, further reinforcing its potential for achieving superior performance.
@@ IV-G1 Kernel Size
To verify the effectiveness of our proposed CCF mechanism, we devise a variant method that substitutes addition for multiplication of $\mathcal{T}_{i,j}$ and $\mathcal{C}_{i,j}$ in the Eq.  6. The results are shown in Tab. LABEL:tab:mva.
TABLE VIII: Test accuracy on three datasets with different CCF operations.
@@ IV-G3 Convergence
We also empirically demonstrate the convergence of our proposed method, as shown in Fig. 9. Specifically, Fig. 9 illustrates the performance trend of vanilla LIF-SNN, Parametric LIF-based SNN [30] without TCJA block and our proposed TCJA-SNN for 1000 epochs. As the training epoch increases, the performance trend of our proposed method becomes more stable and converges to a higher level. Moreover, the TCJA-SNN can achieve the SOTA performance when only training about 260 epochs, which demonstrates the efficacy of the proposed TCJA.
@@ IV-G5 Theoretical Analysis on Receptive Field
Recalling Eq.  4 and Eq.  5, through two 1-D convolutions along different dimensions, we construct two CCS in a vertical relationship, which are stored in $\mathcal{T}$ and $\mathcal{C}$. In particular, TCJA is to construct a CCS, which can perceive a larger area while realizing feature interaction in different directions. This cross-receptive field is able to abolish the limitations caused by the monotonic dimension, thus bringing performance improvements to the network. As a corollary, when the kernel sizes of the two dimensions are the same, we can obtain a square cross-shaped receptive field similar to that of conventional 2-D convolution, which is a effective scheme in 2-D convolution.
@@ IV-G7 Energy Consumption Analysis
Compared to the ANNs, SNNs consumes less energy due to its sparser firing and poorer processing accuracy. Owing to the binary spikes, each operation in SNNs consists of a single floating-point (FP) addition. In ANNs, on the other hand, each operation computes a dot product as a multiply-accumulate (MAC) calculation consisting of one floating-point (FP) multiplication and one FP addition. Consequently, SNNs use less energy than ANNs in the same network design. This discrepancy can also be validated in 45nm CMOS technology, where the energy cost of each SNN operation is 5.1$\times$ lower than that of each 32-bit ANN MAC operation (0.9pJ vs. 4.6pJ) [72], allowing us to examine the energy consumption of each network architecture.
We assess the energy consumption in the network for classifying the DVS128 dataset with both ANN and SNN. first, we assess the spiking rate, floating-point operations per second (FLOPS) of each layer, the result is shown in Tab. IX.
@@
For ANN, we are able to calculate the energy consumption by $FLOPS\times MAC\;energy\;cost$; for SNN, the energy cost should be quantified by $FLOPS\times SNN\;operation\;energy\;cost\times spiking\;rate$. The final power consumption calculation results are $1.90\times 10^{-3}$ (TCJA-SNN) and $10.00\times 10^{-3}$ (ANN), where our TCJA-SNN costs 5.26×5.26$\times$5.26 × lower energy consumption compared to its ANN version.
@@ IV-G8 Propagation Pattern
The forward propagation process in SNNs encompasses both temporal and spatial domains. Intuitively, the computation graph for SNN forward propagation can be conceptualized as a sequential, step-by-step pattern. This pattern is depicted in Fig. 12. In this context, “ step-by-step” refers to the process wherein the network’s output at the initial timestep is evaluated, along with updates to the hidden states of the spiking neurons. Following this, subsequent timesteps are evaluated in a similar manner, maintaining this sequential progression. Besides, the layer-by-layer pattern is also extensively used, which entails performing a spatial forward propagation procedure in which we calculate the output of the first layer at all timesteps as the input of the second layer, then retrieve the output of the last layer at all timesteps. Unlike parallel computing environments like GPU, where the layer-by-layer pattern is preferred, neuromorphic devices operate more like a step-by-step pattern. It can be proved that the output of the network in the two patterns is mathematically equivalent.
@@ V Conclusion
In this paper, we propose the TCJA mechanism, which innovatively recalibrates temporal and channel information in SNN. Specifically, instead of utilizing a generic fully connected network, we use 1-D convolution to build the correlation between frames, reducing the computation and improving model performance. Moreover, we propose a CCF mechanism to realize joint feature interaction between temporal and channel information. Experiments verify the effectiveness of our method with SOTA results on four datasets, i.e., CIFAR10-DVS (83.3%), N-Caltech101 (82.5%), DVS128 (99.0%), Fashion-MNIST (94.8%), CIFAR10 (95.9%) and CIFAR100 (78.3%). In addition to its outstanding performance in classification tasks, TCJA-SNN also exhibits a competitive performance in image generation tasks. To the best of our knowledge, this study represents the pioneering application of the SNN-attention mechanism to both high-level classification and low-level generation tasks. Remarkably, our approach has achieved state-of-the-art performance in both domains, thus making a significant advancement in the field. However, the insertion of TCJA still resulted in a relatively sizable boost in the number of parameters. In future work, we believe this method can easily be integrated into the neuromorphic chip for the hardware-friendly 1-D convolution operation and the binary spiking network structure.